OpenStudy (anonymous):

Use summation notation to write the series 2.2+6.6+11+.. for 5 terms Note: In each of these images, the lower limit of the summation notation is either "n=1" or "n=0".

OpenStudy (anonymous):

\[\sum_{n=0}^{4}2.2 + 4.4n\]

OpenStudy (anonymous):

2.2 + 4.4*0 + 2.2 + 4.4*1 + 2.2 + 4.4*2 + 2.2 + 4.4*3 + 2.2 + 4.4*4 2.2 6.6 11 .......................

OpenStudy (anonymous):

is it -22+4.4n

OpenStudy (anonymous):

\[\sum_{n=0}^{4}2.2\left( 2n+1 \right)\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You can't read what I wrote? My summation expression says 2.2 + 4.4n

OpenStudy (anonymous):

so it is 5 instead of 4

OpenStudy (anonymous):

but that is not one of the choice

OpenStudy (anonymous):

There are many different ways to answer. Didn't know it was multiple choice. My way is one. Surjithayer's is another. Plus there are others

OpenStudy (anonymous):

If they give you answer choices...the easiest way is probably to try each one to see which one give you 2.2 6.6 11 etc

OpenStudy (anonymous):

|

OpenStudy (anonymous):

\[\sum_{n=1}^{5}\left\{ 2.2+\left( n-1 \right)4.4 \right\}\]

OpenStudy (anonymous):

ok that is what I have but it is a negative -2.2

OpenStudy (anonymous):

\[\sum_{n=1}^{5}-2.2+4.4n\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[=\sum_{n=1}^{5}\left\{ 2.2-4.4+4.4n \right\}=\sum_{n=1}^{5}\left( -2.2+4.4 n \right)\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

@surjithayer LOL...How many ways can we think of??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there are so many ways. All are correct.